The result of the nationwide aptitude test in mathematics are normally distributed with a mean of 70 and standard deviation of 5. Find the raw score such that 60% of the cases are below it and case above it.  

1
Expert's answer
2022-03-07T17:13:06-0500

Let X be a random variable representing test result, then X ~ "N(70,5^2)"

Let a be such value that 60% scores are below it, then

"P(X<a)=0.6\implies P(N(70,5^2)<a)=0.6\implies P(70+5N(0,1)<a)=0.6\implies P(N(0,1)<{\frac {a-70} 5})=0.6\implies {\frac {a-70} 5}=0.254\implies a=71.27"

Let b be such value that 60% scores are above it, then, due to symmetric of the normal distribution

"b=70-(a-70)=140-71.27=68.73"